# Graphon Quantum Filtering systems

### v1

@@ 3 Non-exchangeable many-body observed quantum particles
It is not a restriction to proceed in this way, because a complex separable Hilbert space has the following isomorphism $\mathbb{H}\simeq l^{2}_{\mathbb{C}}(\mathbb{N})\simeq L^{2}_{\mathbb{C}}(\mathbb{N},\delta),$ where $\delta$ is a counting measure. For example, if we consider the position of a quantum particle in space, then $\mathfrak{X}=\mathbb{R}^{d}$ with $\mu$ being the Lebesgue measure, if instead we consider discrete energy levels for a quantum particle, we may set $\mathfrak{X}=\mathbb{N}$ or $\mathfrak{X}=\{1,\dots,d\}$, with $\mu$ as the counting measure.
Consider $N$ quantum particles distributed over a graph $G_{N}=(V_{N},E_{N})$, which is completely specified by its adjacency matrix $(\xi_{pq}^{N})_{p,q\in V_{N}}$ indicating the intensity of interaction between two particles, the coefficients $\xi_{pq}^{N}$, issued from the step kernel function $w^{G_{N}}$, can be constructed in two ways:
@@
$\mathbb{H}^{N}:=\bigotimes_{l\in V_{N}}\mathbb{H}$ $=L^{2}_{\mathbb{C}}(\mathfrak{X},\mu)\otimes\dots\otimes L^{2}_{\mathbb{C}}(\mathfrak{X},\mu)\equiv L^{2}_{\mathbb{C}}(\mathfrak{X}^{N},\mu^{\otimes N}),$
@@
$\mathcal{S}(\mathbb{H}^{N})$ :=$:=\Big{\{}\boldsymbol{\rho}\in\mathcal{B}_{1}(\mathbb{H}^{N})\;\Big{|}\;\boldsymbol{\rho}\geq 0,\;\mathrm{t}\mathrm{r}(\boldsymbol{\rho})=1,\;\boldsymbol{\rho}=\boldsymbol{\rho}^{\dagger}\Big{\}}.$
Each particle is governed by a self-adjoint free Hamiltonian $\tilde{H}^{\dagger}=\tilde{H}\in\mathcal{B}(\mathbb{H})$. The interaction between two particles is represented by an operator $A$ on $L^{2}_{\mathbb{C}}(\mathfrak{X}^{2},\mu^{\otimes 2})$ of Hilbert-Schmidt type, with kernel $a$ satisfying the following proprieties:
@@
At the same time, each particle is observed through a quantum channel $L\in\mathcal{B}(\mathbb{H})$ with detection efficiency $\eta\in(0,1]$. Then the dynamics of a posteriori state is given by the $N$-Body Belavkin $\{W^{l},\;l\in V_{N}\}$ of independent Wiener processes
@@
$\quad+\sqrt{\eta}\sum_{q\in V_{N}}\Big{(}{\bf L}_{q}\boldsymbol{\rho}_{t}^{N}+\boldsymbol{\rho}_{t}^{N}{\bf L}^{\dagger}_{q}-\mathrm{t}\mathrm{r}\big{(}({\bf L}_{q}+{\bf L}_{q}^{\dagger})\boldsymbol{\rho}_{t}^{N}\big{)}\boldsymbol{\rho}_{t}^{N}\Big{)}\mathrm{d}W_{t}^{q},$ (3.1)
@@ 4 Mean-field graphon quantum filtering systems
[…]
where $\gamma_{t,u}$ is the density operator of the $u\in I$-th particle at time $t$. Thus, the dynamics of the system in the mean-field limit is given by the following graphon quantum mean-field Belavkin system of equation driven by a familly $\mathbf{W}=\{W^{u}:u\in I\}$ of pairwise independent of Wiener processes on Fubini extended space $(I\times\Omega,\mathcal{V},\mathcal{P})$:
$\mathrm{d}\gamma_{t,{u}}$ $=-\mathrm{i}\Big{[}\tilde{H}+\int_{0}^{1}{w(u,v)}{A^{\mathbb{E}[\gamma_{t,{v}}]}}\mathrm{d}v,\gamma_{t,{u}}\Big{]}\mathrm{d}t+\Big{(}L\gamma_{t,{u}}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,{u}}\}\Big{)}\mathrm{d}t$
$\quad\quad\quad\quad+\sqrt{\eta}\Big{(}L\gamma_{t,{u}}+\gamma_{t,{u}}L^{\dagger}-\mathrm{t}\mathrm{r}\big{(}(L+L^{\dagger})\gamma_{t,{u}}\big{)}\gamma_{t,{u}}\Big{)}\mathrm{d}W_{t}^{u},$ (4.1)
@@
$A^{\rho}(x,y)$ $=\int_{\mathfrak{X}^{2}}a(x,y;x^{\prime},y^{\prime})\overline{\rho(x^{\prime},y^{\prime})}\mu(\mathrm{d}x^{\prime})\mu(\mathrm{d}y^{\prime}),\;\forall\rho\in L^{2}_{\mathbb{C}}(\mathfrak{X}\times\mathfrak{X},\mu^{\otimes 2}).$ (4.2)
@@
An important estimate between the mean-field operator and the pairwise operator is established in [43, Lemma 5.1].
Lemma 1 ([43]).
@@
$\mathrm{d}\gamma_{t,u}$ $=-\mathrm{i}\Big{[}\tilde{H}+A^{\mathbb{E}[\gamma_{t,u}]},\gamma_{t,u}\Big{]}\mathrm{d}t+\Big{(}L\gamma_{t,u}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,u}\}\Big{)}\mathrm{d}t$
@@
In the following, we state our first result, which establishes the wellposedness of the system (4).
Theorem 2 (Wellposedness).
Let $T>0$, $H\in B(\mathbb{H})$ be self-adjoint, $L\in\mathcal{B}_{2}(\mathbb{H})$ and let $A^{\bullet}$ be defined as in (4). Under the Fubini extension probability space $(I\times\Omega,\mathcal{V},\mathcal{P})$ with a family of essentially pairwise independents Wiener processes $\mathbf{W}=\big{\{}W^{u}:u\in I\big{\}}$, the system (4) has a unique strong solution and for every $(t,u)\in[0,T]\times I$, it holds that $\gamma_{t,u}\in\mathcal{S}(\mathbb{H}).$
@@
$\sup_{t\in[0,T]}\mathbb{E}\Big{[}\int_{I}\Big{\|}\tilde{\gamma}_{t,u}-\hat{\gamma}_{t,u}\Big{\|}^{2}_{2}\mathrm{d}u\Big{]}\leq C\|\tilde{w}-\hat{w}\|_{L^{1}}.$
@@
    We formally replace the dependency on law by parameterization and define the mapping $\Xi:\mathcal{C}_{I}\to\mathcal{C}_{I}$ by $\Xi(\xi):=(\Xi_{u}(\xi))_{u\in I},$ where $\Xi_{u}(\xi):=(\mathbb{E}[\gamma_{t,u}^{\xi}])_{0\leq t\leq T}$ with
@@
    $\quad\quad+\sqrt{\eta}\left(L\gamma^{\xi}_{t,u}+\gamma^{\xi}_{t,u}L^{\dagger}-\mathrm{t}\mathrm{r}\left((L+L^{\dagger})\gamma^{\xi}_{t,u}\right)\gamma^{\xi}_{t,u}\right)\mathrm{d}W_{t}^{u},\quad u\in I.$ (4.3)
@@
    $\quad\quad\quad+\sqrt{\eta}\Big{(}L\vartheta^{\xi}_{t,{u}}+\vartheta^{\xi}_{t,{u}}L^{\dagger}\Big{)}\mathrm{d}Y_{t}^{u},\;u\in I,$ (4.4)
@@
    From Proposition 3 in Appendix B, this system admits a strong solution in $\mathcal{L}_{2}(\mathbb{H})$ with positivity preserved for every $u\in I$.
    Define $\gamma_{t,u}^{\xi}:=\frac{\vartheta_{t,u}^{\xi}}{\mathrm{t}\mathrm{r}(\vartheta_{t,u}^{\xi})}.$ By using Proposition (4) and by Itô formula, this corresponds to the solution of Equation (4.3) with
    $W_{t}^{u}:=Y_{t}^{u}-\sqrt{\eta}\int_{0}^{t}\mathrm{t}\mathrm{r}\big{(}(L+L^{\dagger})\gamma_{s,u}^{\xi}\big{)}\mathrm{d}s,$
    where $\mathbf{W}=\big{\{}W^{u}:u\in I\big{\}}$ is a family of essentially pairwise independents Wiener processes on $(I\times\Omega,\mathcal{V},\mathcal{P})$.
@@
    Pathwise uniqueness step: In order to apply the fixed point argument, we need to define a strong solution for any $\xi\in C_{I}$ to (4.3). Since the Liouville space is a Hilbert space, weak existence and pathwise uniqueness imply the existence of a unique strong solution by the Hilbert-space version of the Yamada–Watanabe theorem. Consider two solutions $\gamma^{\xi,a},\gamma^{\xi,b}$ defined on the same probability space $(I\times\Omega,\mathcal{V},\mathcal{P})$ and set $K:=L+L^{\dagger}$. Then,
@@
    $\mathbb{E}\Big{[}\Big{\|}\gamma_{t,u}^{\xi,a}-\gamma_{t,u}^{\xi,b}\Big{\|}_{2}^{2}\Big{]}$ $\leq 4\mathbb{E}\Big{[}t\int_{0}^{t}\Big{\|}\big{[}\tilde{H}+\int_{0}^{1}w(u,v)A^{\xi_{t}^{v}}\mathrm{d}v,(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})\big{]}\Big{\|}_{2}^{2}\mathrm{d}s\Big{]}$
    $+4\mathbb{E}\Big{[}t\int_{0}^{t}$ $\Big{\|}\Big{(}L(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})\}\Big{)}\Big{\|}^{2}_{2}\mathrm{d}s\Big{]}$
    $+4\eta\mathbb{E}\Big{[}\int_{0}^{t}$ $\Big{\|}\Big{(}L{}(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})+(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})L^{\dagger}-\mathrm{t}\mathrm{r}\big{(}K\gamma_{s,u}^{\xi,a}\big{)}(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})\Big{\|}_{2}^{2}\mathrm{d}s\Big{]}$
    $+4\eta\mathbb{E}\Big{[}\int_{0}^{t}$ $\Big{\|}\mathrm{t}\mathrm{r}\big{(}K(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})\big{)}\gamma_{s,u}^{\xi,b}\Big{\|}_{2}^{2}\mathrm{d}s\Big{]}.$
@@
    $\mathrm{t}\mathrm{r}^{2}\big{(}K(\gamma_{t,u}^{\xi,a}-\gamma_{t,u}^{\xi,b})\big{)}$ $\leq\|K\|_{2}\|\gamma_{t,u}^{\xi,a}-\gamma_{t,u}^{\xi,b}\|_{2}^{2}.$
@@
    $\sup_{u\in I}\mathbb{E}\Big{[}\Big{\|}\gamma_{t,u}^{\xi,a}-\gamma_{t,u}^{\xi,b}\Big{\|}_{2}^{2}\Big{]}$ $\leq C\int_{0}^{t}\sup_{u\in I}\mathbb{E}\Big{[}\Big{\|}\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b}\Big{\|}_{2}^{2}\Big{]}\mathrm{d}s.$
    By Grönwall lemma, we conclude $\gamma^{\xi,a}=\gamma^{\xi,b}$ almost surely. Thanks to the Hilbert space formulation of Yamada-Watanabe theorem [62, 30], we obtain the existence of a unique strong solution.
@@
    $\Xi:\mathcal{C}_{I}\to\mathcal{C}_{I}\;\text{by}\;\;\Xi(\xi):=\big{(}(\mathbb{E}_{\mathcal{P}}[\gamma_{t,u}^{\xi}])_{0\leq t\leq T}\big{)}_{u\in I},$
    so that the process $\gamma^{m}$ corresponds to the solution of (4) if and only if $m=\Xi(m)$. Therefore, we should prove existence and uniqueness by showing that the mapping has a unique fixed point, i.e., by proving that $\Xi$ is contraction with respect to the norm on $\mathcal{C}_{I}$.
@@
    $\big{\|}\Xi_{u}(\tilde{\xi})_{t}-\Xi_{u}(\hat{\xi})_{t}$ $\big{\|}_{2}\leq\mathbb{E}\big{[}\|\Delta\gamma_{t,u}\|_{2}\big{]}$
    $\leq\int_{0}^{t}\mathbb{E}\Big{[}\Big{\|}[H,\Delta\gamma_{s,u}]\Big{\|}_{2}+\Big{\|}L\Delta\gamma_{s,u}L^{\dagger}\Big{\|}_{2}+\frac{1}{2}\Big{\|}\{L^{\dagger}L,\Delta\gamma_{s,u}\}\Big{\|}_{2}\Big{]}\mathrm{d}s$
    $+\int_{0}^{t}\Big{(}\mathbb{E}$ $\Big{[}\Big{\|}[\int_{0}^{1}w(u,v)A^{\xi^{v}}\mathrm{d}v,\Delta\gamma_{s,u}]\Big{\|}_{2}\Big{]}+\mathbb{E}\Big{[}\Big{\|}[\int_{0}^{1}w(u,v)\Delta A^{\xi^{v}}\mathrm{d}v,\gamma_{s,{u}}]\Big{\|}_{2}\Big{]}\Big{)}\mathrm{d}s$
    $+\int_{0}^{t}\Big{(}\mathbb{E}$ $\Big{[}\Big{\|}L\Delta\gamma_{s,u}+\Delta\gamma_{s,u}L^{\dagger}\Big{\|}_{2}^{2}\Big{]}\Big{)}^{\frac{1}{2}}\mathrm{d}s$
    $+\int_{0}^{t}\Big{(}\mathbb{E}\Big{[}\Big{\|}\mathrm{t}\mathrm{r}\big{(}K\Delta\gamma_{s,u}\big{)}\gamma_{s,u}^{\tilde{\xi}}\Big{\|}_{2}^{2}\Big{]}\Big{)}^{\frac{1}{2}}+\Big{(}\mathbb{E}\Big{[}\Big{\|}\mathrm{t}\mathrm{r}\big{(}K\gamma_{s,u}^{\hat{\xi}}\big{)}\Delta\gamma_{s,u}\Big{\|}_{2}^{2}\Big{]}\Big{)}^{\frac{1}{2}}\mathrm{d}s$
    $\leq C\int_{0}^{t}\Big{(}\mathbb{E}\Big{[}\Big{\|}\Delta\gamma_{s,u}\Big{\|}_{2}\Big{]}+\int_{0}^{1}w(u,v)\big{\|}\Delta\xi_{s}^{v}\big{\|}_{2}\mathrm{d}v\Big{)}\mathrm{d}s.$

### v3

@@ 3 Non-exchangeable many-body observed quantum particles
It is not a restriction to proceed in this way, because a complex separable Hilbert space has the following isomorphism $\mathbb{H}\simeq l^{2}_{\mathbb{C}}(\mathbb{N})\simeq L^{2}_{\mathbb{C}}(\mathbb{N},\delta),$ where $\delta$ is a counting measure. For example, if we consider the position of a quantum particle in space, then $\mathfrak{X}=\mathbb{R}^{d}$ with $\mu$ being the Lebesgue measure. If instead we consider discrete energy levels for a quantum particle, we may set $\mathfrak{X}=\mathbb{N}$ or $\mathfrak{X}=\{1,\dots,d\}$, with $\mu$ as the counting measure.
Consider $N$ quantum particles distributed over a graph $G_{N}=(V_{N},E_{N})$, which is completely specified by its adjacency matrix $(\xi_{pq}^{N})_{p,q\in V_{N}}$ indicating the intensity of interaction between two particles. The coefficients $\xi_{pq}^{N}$, issued from the step kernel function $w^{G_{N}}$, can be constructed in two ways:
@@
$\mathbb{H}^{N}:=\bigotimes_{q\in V_{N}}\mathbb{H}$ $=L^{2}_{\mathbb{C}}(\mathfrak{X},\mu)\otimes\dots\otimes L^{2}_{\mathbb{C}}(\mathfrak{X},\mu)\equiv L^{2}_{\mathbb{C}}(\mathfrak{X}^{N},\mu^{\otimes N}),$
@@
$\mathcal{S}(\mathbb{H}^{N})$ $:=\Big\{\boldsymbol{\rho}\in\mathcal{B}_{1}(\mathbb{H}^{N})\;\Big|\;\boldsymbol{\rho}\geq 0,\;\mathrm{t}\mathrm{r}(\boldsymbol{\rho})=1,\;\boldsymbol{\rho}=\boldsymbol{\rho}^{\dagger}\Big\}.$
[…]
@@
At the same time, each particle is observed through a quantum channel $L\in\mathcal{B}(\mathbb{H})$ with detection efficiency $\eta\in(0,1]$. Then the dynamics of a posteriori state is given by the $N$-Body Belavkin, with a family $\mathbf{W}^{N}=\{W^{q},\;q\in V_{N}\}$ of pairwise independent Wiener processes
@@
$\quad+\sqrt{\eta}\sum_{q\in V_{N}}\Big({\bf L}_{q}\boldsymbol{\rho}_{t}^{N}+\boldsymbol{\rho}_{t}^{N}{\bf L}^{\dagger}_{q}-\mathrm{t}\mathrm{r}\big(({\bf L}_{q}+{\bf L}_{q}^{\dagger})\boldsymbol{\rho}_{t}^{N}\big)\boldsymbol{\rho}_{t}^{N}\Big)\mathrm{d}W_{t}^{q},$ (5)
@@ 4 Mean-field graphon quantum filtering systems
$\boldsymbol{\rho}_{0}^{N}=\rho_{0}^{\otimes N}\xrightarrow[N\to\infty]{\text{Propagation of chaos}}\boldsymbol{\rho}_{t}^{N}\approx\bigotimes_{u\in I}\gamma_{t,u},$
where $\gamma_{t,u}$ is the density operator of the $u\in I$-th particle at time $t$.
Thus, the dynamics of the system in the mean-field limit are given by the following graphon quantum mean-field Belavkin system of equation driven by a familly $\mathbf{W}=\{W^{u}:u\in I\}$ of e.p.i. Wiener processes on Fubini extended space $(I\times\Omega,\mathcal{V},\mathcal{P})$ of the usual product space $(I\times\Omega,\mathcal{I}\otimes\mathcal{F},\lambda\otimes\mathbb{P})$:
$\mathrm{d}\gamma_{t,{u}}$ $=-\mathrm{i}\Big[\tilde{H}+\int_{0}^{1}{w(u,v)}{A^{\mathbb{E}_{\mathbb{P}}[\gamma_{t,{v}}]}}\mathrm{d}v,\gamma_{t,{u}}\Big]\mathrm{d}t+\Big(L\gamma_{t,{u}}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,{u}}\}\Big)\mathrm{d}t$
$\quad\quad\quad\quad+\sqrt{\eta}\Big(L\gamma_{t,{u}}+\gamma_{t,{u}}L^{\dagger}-\mathrm{t}\mathrm{r}\big((L+L^{\dagger})\gamma_{t,{u}}\big)\gamma_{t,{u}}\Big)\mathrm{d}W_{t}^{u},$ (6)
@@
$A^{\rho}(x,x^{\prime})$ $=\int_{\mathfrak{X}^{2}}a(x,y;x^{\prime},y^{\prime})\overline{\rho(y,y^{\prime})}\mu(\mathrm{d}y)\mu(\mathrm{d}y^{\prime}),\;\forall\rho\in L^{2}_{\mathbb{C}}(\mathfrak{X}\times\mathfrak{X},\mu^{\otimes 2}).$ (7)
@@
An important estimate between the mean-field operator and the pairwise operator is established in [47, Lemma 5.1].
Lemma 4.1 ([47]).
@@
$\mathrm{d}\gamma_{t,u}$ $=-\mathrm{i}\Big[\tilde{H}+A^{\mathbb{E}_{\mathbb{P}}[\gamma_{t,u}]},\gamma_{t,u}\Big]\mathrm{d}t+\Big(L\gamma_{t,u}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,u}\}\Big)\mathrm{d}t$
@@
Remark 4.3 (Formulation in the Fubini extended space).
Before stating the first theorem, it is useful to point out that the graphon SDEs (4) can be formulated as a single SDE defined on the Fubini extension space, driven by a standard Wiener process (see [2] for details and [30] for an $L^{2}$–formulation approach).
Let
$\mathcal{H}:=L^{2}_{\mathcal{L}_{2}(\mathbb{H})}(I\times\Omega,\mathcal{P}),\qquad\mathcal{H}_{I}:=L^{2}_{\mathcal{L}_{2}(\mathbb{H})}(I,\lambda).$
The collection $\{\gamma_{t,u}:u\in I\}$ can equivalently be represented by
$\Gamma_{t}\in\mathcal{H},\qquad\Gamma_{t}(u,\omega):=\gamma_{t,u}(\omega).$
Define
$\mathrm{M}_{t}:=\mathbb{E}_{\mathbb{P}}[\Gamma_{t}]\in\mathcal{H}_{I},$
so that $\mathrm{M}_{t}(u)=\mathbb{E}_{\mathbb{P}}[\gamma_{t,u}]$ for each $u\in I$.
Given the Hilbert–Schmidt interaction operator $A^{\bullet}$, the linear operator of interaction $\mathbb{W}_{A}$ is defined by
$\mathbb{W}_{A}:$ $\;\mathcal{H}_{I}\longrightarrow\mathcal{H}_{I}$
$\big(\mathbb{W}_{A}[\mathrm{G}]\big)(u)$ $:=\int_{I}w(u,v)A^{\mathrm{G}(v)}\mathrm{d}v,\qquad\forall\mathrm{G}\in\mathcal{H}_{I},\;u\in I.$
With these notations, the system (4) can be written as the Hilbert-space valued stochastic equation
$\mathrm{d}\Gamma_{t}=\mathcal{D}\big[\Gamma_{t},\mathbb{W}_{A}(\mathrm{M}_{t})\big]\mathrm{d}t+\mathcal{K}[\Gamma_{t}]\mathrm{d}\mathbf{W}_{t},$ (8)
where:
  • $\mathcal{D}[.,.]$ collects the drift terms:
    $\mathcal{D}[\Gamma,\Gamma^{\prime}]=-\mathrm{i}[\tilde{H}+\Gamma^{\prime},\Gamma]+L\Gamma L^{\dagger}-\tfrac{1}{2}\{L^{\dagger}L,\Gamma\},$
  • $\mathcal{K}[.]$ denotes the diffusion part:
    $\mathcal{K}[\Gamma]=\sqrt{\eta}\,\big(L\Gamma+\Gamma L^{\dagger}-\mathrm{t}\mathrm{r}((L+L^{\dagger})\Gamma)\Gamma\big).$
The following theorem establishes the well-posedness of the system (4).
Theorem 4.4 (Well-posedness).
[…]
@@
$\sup_{t\in[0,T]}\mathbb{E}_{\mathbb{P}}\Big[\int_{I}\Big\|\tilde{\gamma}_{t,u}-\hat{\gamma}_{t,u}\Big\|^{2}_{2}\mathrm{d}u\Big]\leq C\|\tilde{w}-\hat{w}\|_{L^{1}}.$
@@
    We formally replace the dependency on law by parameterization and define the mapping $\Xi:\mathcal{C}_{I}\to\mathcal{C}_{I}$ by $\Xi(\xi):=(\Xi_{u}(\xi))_{u\in I},$ where $\Xi_{u}(\xi):=(\mathbb{E}_{\mathbb{P}}[\gamma_{t,u}^{\xi}])_{0\leq t\leq T}$ with
@@
    $\quad\quad+\sqrt{\eta}\left(L\gamma^{\xi}_{t,u}+\gamma^{\xi}_{t,u}L^{\dagger}-\mathrm{t}\mathrm{r}\left((L+L^{\dagger})\gamma^{\xi}_{t,u}\right)\gamma^{\xi}_{t,u}\right)\mathrm{d}W_{t}^{u},\quad u\in I.$ (9)
@@
    $\quad\quad\quad+\sqrt{\eta}\Big(L\vartheta^{\xi}_{t,{u}}+\vartheta^{\xi}_{t,{u}}L^{\dagger}\Big)\mathrm{d}Y_{t}^{u},\;u\in I,$ (10)
@@
    From Proposition B.2 in Appendix B, this system admits a strong solution in $\mathcal{L}_{2}(\mathbb{H})$ with positivity preserved for every $u\in I$.
    Define $\gamma_{t,u}^{\xi}:=\frac{\vartheta_{t,u}^{\xi}}{\mathrm{t}\mathrm{r}(\vartheta_{t,u}^{\xi})}.$ By using Proposition (B.3) and by Itô formula, this corresponds to a weak solution of Equation (9). Note that $\gamma_{t,u}^{\xi}$ is positive and trace-one preserving as soon as $\mathrm{t}\mathrm{r}(\gamma_{0,u}^{\xi})=1$.
@@
    Pathwise uniqueness step: In order to apply the fixed point argument, we need to define a strong solution for any $\xi\in\mathcal{C}_{I}$ to (9). Since the Liouville space is a Hilbert space, weak existence and pathwise uniqueness imply the existence of a unique strong solution by the Hilbert-space version of the Yamada–Watanabe theorem. Consider two solutions $\gamma^{\xi,a},\gamma^{\xi,b}$ defined on the same probability space $(I\times\Omega,\mathcal{V},\mathcal{P})$ with the same initial state in $\mathcal{L}_{2}(\mathbb{H})$, and set $K:=L+L^{\dagger}$. Then,
@@
    $\mathbb{E}_{\mathbb{P}}\Big[\Big\|\gamma_{t,u}^{\xi,a}-\gamma_{t,u}^{\xi,b}\Big\|_{2}^{2}\Big]$ $\leq 4\mathbb{E}_{\mathbb{P}}\Big[t\int_{0}^{t}\Big\|\big[\tilde{H}+\int_{0}^{1}w(u,v)A^{\xi_{t}^{v}}\mathrm{d}v,(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})\big]\Big\|_{2}^{2}\mathrm{d}s\Big]$
    $+4\mathbb{E}_{\mathbb{P}}\Big[t\int_{0}^{t}$ $\Big\|\Big(L(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})\}\Big)\Big\|^{2}_{2}\mathrm{d}s\Big]$
    $+4\eta\mathbb{E}_{\mathbb{P}}\Big[\int_{0}^{t}$ $\Big\|\Big(L{}(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})+(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})L^{\dagger}-\mathrm{t}\mathrm{r}\big(K\gamma_{s,u}^{\xi,a}\big)(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})\Big\|_{2}^{2}\mathrm{d}s\Big]$
    $+4\eta\mathbb{E}_{\mathbb{P}}\Big[\int_{0}^{t}$ $\Big\|\mathrm{t}\mathrm{r}\big(K(\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b})\big)\gamma_{s,u}^{\xi,b}\Big\|_{2}^{2}\mathrm{d}s\Big].$
@@
    $\mathrm{t}\mathrm{r}^{2}\big(K(\gamma_{t,u}^{\xi,a}-\gamma_{t,u}^{\xi,b})\big)$ $\leq\|K\|_{2}^{2}\|\gamma_{t,u}^{\xi,a}-\gamma_{t,u}^{\xi,b}\|_{2}^{2}.$
@@
    $\sup_{u\in I}\mathbb{E}_{\mathbb{P}}\Big[\Big\|\gamma_{t,u}^{\xi,a}-\gamma_{t,u}^{\xi,b}\Big\|_{2}^{2}\Big]$ $\leq C\int_{0}^{t}\sup_{u\in I}\mathbb{E}_{\mathbb{P}}\Big[\Big\|\gamma_{s,u}^{\xi,a}-\gamma_{s,u}^{\xi,b}\Big\|_{2}^{2}\Big]\mathrm{d}s.$
    By Grönwall lemma, we conclude $\gamma^{\xi,a}=\gamma^{\xi,b}$ almost surely. Thanks to the Hilbert space formulation of Yamada-Watanabe theorem (see [53, Appendix E]), we obtain the existence of a unique strong solution.
@@
    $\Xi:\mathcal{C}_{I}\to\mathcal{C}_{I}\;\text{by}\;\;\Xi(\xi):=\big((\mathbb{E}_{\mathbb{P}}[\gamma_{t,u}^{\xi}])_{0\leq t\leq T}\big)_{u\in I},$
    so that the process $\gamma^{m}$ corresponds to the solution of (4) if and only if $m=\Xi(m)$. Therefore, we should prove existence of a unique fixed point, i.e., by proving that $\Xi$ is contraction with respect to the norm on $\mathcal{C}_{I}$.
@@
    $\big\|\Xi_{u}(\tilde{\xi})_{t}-\Xi_{u}(\hat{\xi})_{t}$ $\big\|_{2}\leq\mathbb{E}_{\mathbb{P}}\big[\|\Delta\gamma_{t,u}\|_{2}\big]$
    $\leq\int_{0}^{t}\mathbb{E}_{\mathbb{P}}\Big[\Big\|[H,\Delta\gamma_{s,u}]\Big\|_{2}+\Big\|L\Delta\gamma_{s,u}L^{\dagger}\Big\|_{2}+\frac{1}{2}\Big\|\{L^{\dagger}L,\Delta\gamma_{s,u}\}\Big\|_{2}\Big]\mathrm{d}s$
    $+\int_{0}^{t}\Big(\mathbb{E}_{\mathbb{P}}$ $\Big[\Big\|[\int_{0}^{1}w(u,v)A^{\xi^{v}}\mathrm{d}v,\Delta\gamma_{s,u}]\Big\|_{2}\Big]+\mathbb{E}_{\mathbb{P}}\Big[\Big\|[\int_{0}^{1}w(u,v)\Delta A^{\xi^{v}}\mathrm{d}v,\gamma_{s,{u}}]\Big\|_{2}\Big]\Big)\mathrm{d}s$
    $+\int_{0}^{t}\Big(\mathbb{E}_{\mathbb{P}}$ $\Big[\Big\|L\Delta\gamma_{s,u}+\Delta\gamma_{s,u}L^{\dagger}\Big\|_{2}^{2}\Big]\Big)^{\frac{1}{2}}\mathrm{d}s$
    $+\int_{0}^{t}\Big(\mathbb{E}_{\mathbb{P}}\Big[\Big\|\mathrm{t}\mathrm{r}\big(K\Delta\gamma_{s,u}\big)\gamma_{s,u}^{\tilde{\xi}}\Big\|_{2}^{2}\Big]\Big)^{\frac{1}{2}}+\Big(\mathbb{E}_{\mathbb{P}}\Big[\Big\|\mathrm{t}\mathrm{r}\big(K\gamma_{s,u}^{\hat{\xi}}\big)\Delta\gamma_{s,u}\Big\|_{2}^{2}\Big]\Big)^{\frac{1}{2}}\mathrm{d}s$
    $\leq C\int_{0}^{t}\Big(\mathbb{E}_{\mathbb{P}}\Big[\Big\|\Delta\gamma_{s,u}\Big\|_{2}\Big]+\int_{0}^{1}w(u,v)\big\|\Delta\xi_{s}^{v}\big\|_{2}\mathrm{d}v\Big)\mathrm{d}s.$